OpenStudy (anonymous):

What’s the equation of a line that passes through points (0, -1) and (2, 3)?

OpenStudy (e.mccormick):

OK, so where did you get stuck on this?

OpenStudy (e.mccormick):

This is the line it is talking about. The goal is to find an equation that defines it.

OpenStudy (e.mccormick):

Usually you want to do this by putting it in the form of y=mx+b. You have a y and x, two in fact. So what you need to find is m, the slope, and b, the y-intercept. This one was very nice because they even gave you the y intercept.

OpenStudy (anonymous):

apply the (two point form) formula.you get {(y+1)/4}=(x/2).

OpenStudy (anonymous):

You are given two coordinates. Use those two to find the slope or the "m" in: \[y=mx+b\]. Keep in mind which is is which so you get the correct slope. (0, -1) and (2, 3). Use the formula below to find the slope: \[\frac{ Y_{2}-Y_{1} }{ X_{2}-X_{1} }\] Now plugin the values you are given: \[\frac{ 3-(-1) }{ 2-0 }\] Which becomes: \[\frac{ 3+1 }{ 2-0 }\] So your slope is: \[\frac{ 4 }{ 2 }\] or simply 2. Now that you have your "m" for y=mx+b, you now have y=2x+B To find what B is, use one of the coordinates. We'll use (0,-1) in this case. y=2x+B -1=2(0)+B -1=B. There's your B, you now have found the whole equation. y=2x-1. :)

OpenStudy (e.mccormick):

@ch3n2x Yep. I was hoping whenever they came back the clues I left would be enough. Don't even know how far they got on it.

OpenStudy (anonymous):

(2x-y-1=0) is the equation.